# Calibrating Long-form Generations from Large Language Models

### v1

@@ 6.4 Application
We showcase a cost-effective application of our system (illustrated in Figure 5) by using a cascade of language models with varying capabilities for selectively responding to queries, aiming for optimal performance within a limited API budget. We initially employ a set of calibrated open-source LMs (selected by our calibration framework) to handle queries. Complex queries, identified by the lower confidence of these models (derived from our self-consistency confidence elicitation method), are escalated to more advanced API LMs. This tiered approach strikes a balance between cost-efficiency and performance in LLM applications.
@@
Our experiments, conducted on the ASQA and QAMPARI datasets, involve cascading a series of calibrated open-source models, culminating with GPT-3.5-turbo. See Appendix D.3 for model selection and cascading details. We begin with the first open-source model, addressing questions where it believes the answer has a probability higher than $\tau_{c}$ that the answer’s correctness score is higher than $\tau_{s}$=0.6. Subsequent questions are then processed through the remaining open-source models with the same method. Any queries not confidently handled by these models are ultimately directed to GPT-3.5-turbo for resolution. By adjusting $\tau_{c}$ from 0 to 1, we can vary the number of queries that will be sent to GPT-3.5-turbo, allowing us to analyze the quality of the final answer as the budget increases for calling the API. We use the strongest open-source model’s results as a baseline for a zero API budget, and GPT-3.5-turbo’s results for a full budget scenario. As a baseline, we compare to a strategy where, within a limited number of API requests, a random selection of queries is addressed by GPT-3.5-turbo and the rest are handled by the strongest open-source model. For each API budget, we randomly select questions for GPT-3.5-turbo using 10 different random seeds and calculate the mean and std for the results. As Figure 6 shows, our LLM cascade method always outperforms the random response strategy across various budget levels. We don’t consider the API cost for obtaining self-consistency confidence scores from open-source models, since GPT-3.5 could be substituted with other open-source NLI models for verifying sentence presence in the second answer.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111606060606565656570707070BudgetCorrectnessASQA
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111151515152020202025252525BudgetCorrectnessQAMPARI
Figure 6: Variation in Expected Correctness Score with API Budget Allocation on ASQA and QAMPARI: This graph contrasts the LLM Cascade approach (blue) with the Random Response strategy (red). The red line shows the mean correctness score from 10 random seeds, while the light red shaded area indicates the range of standard deviation. Notably, a budget of 1 corresponds to the correctness achieved by GPT-3.5-turbo, while a budget of 0 reflects the performance of the best-performing open-source model selected for LLM cascade.
@@ 7 Conclusion
In conclusion, our research introduces a comprehensive calibration system designed to evaluate LLMs across both short-form and long-form text generation. We have developed confidence elicitation methods that derive confidence distributions from LLMs. To assess calibration, we introduced three new metrics, each tailored to capture different aspects of LLM performance. Our findings reveal that larger LLMs do not necessarily equate to better calibration, and the degree of calibration significantly varies across different datasets. This challenges prevailing assumptions in the field and underscores the need for nuanced approaches in LLM development and assessment. Furthermore, our study explores various strategies to enhance LLM calibration, offering practical solutions to common issues faced in the field. We also present a potential application of our system, demonstrating its effectiveness in optimizing LLMs performance within the constraints of limited API budgets. The system and tools we present are crucial for further improving the liability of LLMs.
Our research highlights a couple of research directions, including (1) Advancing Calibration Techniques. Develop more sophisticated methods like integrating human feedback into the calibration process, incorporating machine learning algorithms that can adaptively adjust calibration based on the context or type of query, and so on. (2) Cross-Domain Applications. Extend the application of our system to other domains such as medical text generation, legal document analysis, education chatbots and so on. (3) Real-time Calibration. Develop methods for real-time calibration of LLMs in dynamic environments (e.g. interactive chatbots, updated knowledge base). (4) Pre-training with Calibration. Incorporate calibration into the pre-training phase of LLMs. (5) Ethical and Biases Consideration. Investigating how different calibration strategies affect the fairness and bias in LLM outputs can be crucial for responsible AI. (6) Scalability and Efficiency. Explore ways to improve the scalability and efficiency of our calibration system, making it applicable to large-scale, real-world scenarios.
@@ Appendix A Dataset
CNNDM (Nallapati et al., 2016) is a large-scale news summarization dataset containing news articles from CNN [^1] https://www.cnn.com/ and DailyMail [^2] https://www.dailymail.co.uk/. The original CNNDM dataset consists of both source news articles and reference summaries. However, recent work Liu et al. (2023b); Zhang et al. (2023a) has found that the provided reference summaries are not of very good quality and zero-shot LLMs summaries are preferred by human annotators over the reference summaries.
@@ Appendix B Self-consistency
Naive The most basic approach utilizes an additional LLM (GPT-3.5-turbo in our experiments, which can be replaced by other models trained for this task) to determine if two responses are akin, assigning a corresponding similarity score. This method diverges from the relevant technique in contemporary research (Chen et al., 2023b), which primarily focuses on identifying the most consistent answer. Instead, our approach aims to secure specific consistency ratings that reflect the model’s assurance in its primary answer, offering a general overview of the answers’ similarity.
Sentence Split For a more detailed similarity analysis between the two answers, we split the first answer into individual sentences. Another LLM (GPT-3.5-turbo or a similar NLI model) is then used to evaluate whether similar statements are present in the second answer. This method’s limitation is that not all sentences carry equal informational weight. Some may be filler or less informative, potentially skewing the similarity assessment.
@@
Appendix C Main Experiment
@@ C.1 Generation
C.2 Correctness Evaluation
For ASQA, ELI5, CNNDM, we ask the GPT-4 to evaluate the correctness of examples. For ASQA and ELI5, We ask GPT-4 to evaluate the answer three times, producing a distribution capturing criteria ambiguity and model’s inherent subjectivity. However, for CNNDM, we only ask GPT-4 to evaluate once to save computation given that most answers correctness concentrate around 0.8 with small variance. The GPT-4 evaluation template can be found in Appendix E.1 and criteria for different tasks can be found in Appendix E.2. For QAMPARI, which involves generating a list of entities as answers, we determine that F1 scores provide a more suitable measure of correctness than GPT-4 evaluations. To this end, we adopt the F1-5 metric (Gao et al., 2023), computing the F1 score based on direct matches with the correct answer list and assigning a recall of 100% for responses containing at least five accurate answers (recall-5).
C.3 Confidence Elicitation
@@
Self-Evaluation We prompt the model to self-evaluate an answer 10 times, creating a confidence distribution based on its self-evaluation scores. Each self-evaluation includes task instructions, grading criteria, and evaluation examples (three for all long-form QA datasets and one for CNNDM due to length constraints). Additionally, it contains specific instructions for self-evaluation, incorporating both the question and the answer under evaluation. The detailed self-evaluation template and criteria are available in Appendix E.3 and Appendix E.2 respectively.
@@
For ASQA, we employ a self-consistency-claim approach. To assess the similarity between the two answers, we first identify factual claims in the first answer using a ClaimBuster-trained DeBERTa V2 (He et al., 2020) claim detector. We then verify the presence of these claims in the second answer using GPT-3.5-turbo. The similarity score is the average presence of factual claims across answers.
@@
For ELI5, where answers provide easily understandable explanations, we focus on the overall content. Here, we apply a self-consistency-naive method, asking an additional model (GPT-3.5-turbo) to assign a similarity score to each answer pair.
@@
In CNNDM, where answers are summaries highlighting key points of an article, we gauge the similarity between two answers by evaluating the overlap of key points. To achieve this, we implement the self-consistency-split method. We dissect the first answer into individual sentences and then use GPT-3.5-turbo to determine if each sentence is present in the second answer. The similarity score is derived by averaging the presence of these segmented sentences in the comparative answer.
Appendix D Improving Calibration
D.1 Fine-tuning
@@
In our study, we focus on improving model calibration on the ASQA dataset. We generate a self-evaluation training dataset with GPT-4, which evaluates different models responses to questions drawn from ASQA training set. We create the self-evaluation data in two steps. The initial phase involves the creation of a diverse answer pool. This is achieved by employing a suite of models with varying computational capacities, including Llama-2-7b, Llama-2-13b, Llama2-70b, Vicuna-13b, and ChatGPT. Each model generates responses to a spectrum of questions drawn from the training set of ASQA task, ensuring the resultant answer pool encompasses a broad quality spectrum, from low to high. Subsequently, we employ GPT-4 to critically assess these answers, assigning a score and providing a corresponding explanation for its evaluation. This process yields a rich dataset, each instance of which encompasses a question, a model-generated answer, an evaluative score, and a justification for that scoring. This approach results in a comprehensive dataset with 2,000 self-evaluation examples (1,800 for training and 200 for validation), each including a question, model-generated answer, score, and justification. For generation data, we use 80% of ASQA’s dataset (4,353 examples) for training and the remaining 20% for validation. The hybrid dataset combines the self-evaluation and generation training sets, using the self-evaluation validation set for assessment.
@@
Regarding training, we fine-tune Llama2-13b-chat model using LORA on this dataset. We maintain consistent parameters across all scenarios: a learning rate of 5e-6, five epochs, 100 warm-up steps, and a total batch size of 4 (achieved through 4 gradient accumulation steps across four GPUs, with a batch size of 1 per device).
D.2 Hybrid Confidence Elicitation
We integrate self-evaluation and self-consistency methods to assess their combined impact on model calibration. Experiments are performed on the ASQA dataset using Llama-2-13b, Llama-2-70b, and Vicuna-13b models, adjusting the self-consistency weight from 0 to 1. Our findings indicate that this hybrid approach improves correlation, though it does not significantly affect other metrics.
[…]
@@ D.2 Hybrid Confidence Elicitation
D.3 LLM Cascade
Our research utilizes ASQA and QAMPARI datasets, applying self-consistency confidence in our methodology. We use the initial 500 samples from the evaluation data provided in Gao et al. (2023) as our test set and samples from 700 to 800 as the validation set. Initially, we assess the calibration of models Llama-2-13b, Llama-2-70b, and Llama-2-70b on the validation set, with the findings presented in Table 4. Notably, the Llama-2-70b model exhibits inferior calibration compared to the other models when tested on QAMPARI, leading us to exclude it from the model cascade for the test set. In the cascading process, we prioritize models based on their performance, ensuring that the most capable model addresses questions that multiple models confidently identify. Consequently, the sequence of cascaded models for ASQA is Llama-2-70b, Llama-2-13b, Vicuna-13b, GPT-3.5-turbo, and for QAMPARI, it is Llama-2-13b, Vicuna-13b, GPT-3.5-turbo. Then we perform our pipeline on the test sets for these tasks.
[…]
ASQA
Vicuna-13b 50.3 70.7 49.2 47.4
Llama-2-13b 51.7 76.4 21.3 47.6
Llama-2-70b 48.0 76.0 45.2 60.4
QAMPARI
Vicuna-13b 48.8 75.5 27.3 13.5
Llama-2-13b 43.0 66.2 45.7 14.1
Llama-2-70b 3.4 56.0 20.0 8.9
Table 4: Open-source models results on validation sets of ASQA and QAMPARI
Appendix E Prompts

### v2

@@ 4.5 Application
We showcase an application of long-form calibration in Figure 4: a cost-effective cascading strategy using language models of varying capabilities to efficiently handle queries within an API budget constraint. Initially, an open-source model (Llama-2-13b in our experiment) address questions where it believes the answer has a probability higher than $\tau_{c}$ that the answer’s correctness score is above $\tau_{s}$. Complex queries, flagged by lower model confidence, are escalated to a more advanced API LM (GPT-4). Adjusting $\tau_{c}$ between 0 and 1 controls how many queries reach GPT-4, balancing answer quality with API budget constraints. We benchmark using the open-source LM for a zero API budget and the commercial LM for full-budget scenarios.
@@
Our experiments utilize the ASQA and QAMPARI datasets to evaluate four distinct confidence elicitation strategies: PSC, ASC, CSE, and BSE. Additionally, we incorporate a baseline strategy where, under a constrained number of API requests, a random selection of queries is processed by Llama-2-13b, with the remaining handled by GPT-4. For each API budget scenario, questions are randomly assigned to Llama-2-13b using 10 different random seeds, and we calculate the mean and standard deviation of the results. We focus on the success rate, which we define as the percentage of answers that meet or exceed a user-specified score threshold. This metric is reported both for the overall dataset and for the subset of queries selected and handled by Llama-2-13b, illustrating both the general effectiveness of our cascading model and the selective answering capabilities of the individual model. As shown in Figure 5, PSC generally outperforms the other methods, with CSE and BSE yielding comparable results that follow. ASC, in contrast, performs the poorest, comparable to the random selection strategy. These results highlight the pivotal role of advanced calibration techniques, confidence elicitation methods in our case, in boosting the practical utility and cost-efficiency of LLMs when API usage is limited.
(a) ASQA Overall
(b) ASQA Selected
(c) QAMPARI Overall
(d) QAMPARI Selected
Figure 5: Variation in Success Rate by API Budget Allocation on the ASQA and QAMPARI Datasets for All Queries and Those Selected by Llama-2.
@@ 5 Conclusion
Our study presents a novel calibration system for evaluating LLMs in long-form generation. Our results challenge the assumption that larger LLMs are always calibrated better and show calibration variability across datasets. Additionally, we propose methods to improve LLM calibration and show an application that optimizes performance under API budget constraints. The system we present is crucial for further improving the liability of LLMs.
6 Limitation
Our study faces three primary limitations. First, we rely on GPT-4 to estimate the target correctness distribution. But as tasks become more subjective, consensus on humans’ evaluations may decrease. This wider target correctness distribution becomes challenging for both GPT-4 and human annotators to accurately capture. This limitation is inherent to natural language generation (NLG) evaluation and lies beyond the purview of our project. Our framework operates under the premise that a target correctness distribution exists and concentrates on calibration which aligns the model’s confidence with this assumed target. Second, our experiments focus on long-form QA and do not extend to specialized domains such as law, medicine, or education, where the calibration of LLMs could have significant real-world implications. Lastly, our self-consistency method is computationally intensive, posing a challenge for practical applications. There is a need for more efficient approaches in real-world settings.
Replicability:
Codes: https://github.com/kkkevinkkkkk/calibration
@@ Appendix A Dataset
CNNDM (Nallapati et al., 2016) is a large-scale news summarization dataset containing news articles from CNN [^1] https://www.cnn.com/ and DailyMail [^2] https://www.dailymail.co.uk/. The original CNNDM dataset consists of both source news articles and reference summaries. However, recent work Liu et al. (2023b); Zhang et al. (2023) has found that the provided reference summaries are not of very good quality and zero-shot LLMs summaries are preferred by human annotators over the reference summaries.
@@ Appendix B Self-consistency
Naive The most basic approach utilizes an additional LLM (GPT-3.5-turbo in our experiments, which can be replaced by other models trained for this task) to determine if two responses are akin, assigning a corresponding similarity score. This method diverges from the relevant technique in contemporary research (Chen et al., 2023b), which primarily focuses on identifying the most consistent answer. Instead, our approach aims to secure specific consistency ratings that reflect the model’s assurance in its primary answer, offering a general overview of the answers’ similarity. The template for similarity measuring can be found in Appendix F.4.
Sentence Split For a more detailed similarity analysis between the two answers, we split the first answer into individual sentences. Another LLM (GPT-3.5-turbo or a similar NLI model) is then used to evaluate whether similar statements are present in the second answer. This method’s limitation is that not all sentences carry equal informational weight. Some may be filler or less informative, potentially skewing the similarity assessment. The template can be found in Appendix F.4.
@@
Appendix C Evaluation Metric
C.1 GPT-4 metric
We ask GPT-4 to range an answer from 0 (worst) to 5 (best), which is then normalized to [0, 1]. See Appendix F.1 for details of the template. To mitigate scoring variability from criteria ambiguity and LLM uncertainty, we have evaluators repeatedly score each answer, forming a score distribution that better reflects its correctness. Specifically, given an LLM evaluator $\mathcal{E}$, $N$ evaluations from it, and a score $x\in[0,1]$, the correctness distribution is:
$P_{T_{i}}(x)$ $=\frac{1}{N}\sum_{j=1}^{N}\mathrm{1}(\mathcal{E}(A_{i})_{j}=x)$ (12)
C.2 Human Evaluation
We utilize GPT-4 to assess answers across different tasks, including ASQA, ELI5, and CNNDM, using GPT-4 scores as a proxy for the target distribution. To demonstrate the better alignment of GPT-4 scores with human preferences over task-specific metrics, we focus our human evaluation efforts on the long-form tasks of ASQA and ELI5. This approach is supported by prior research indicating GPT-4’s congruence with human judgments on summarization tasks (Liu et al., 2023a), thereby obviating the need for manual evaluation of CNNDM. Following Gao et al. (2023), in ASQA, we adopt the EM-recall automatic metric, which gauges the recall of correct short answers by verifying if the dataset’s provided short answers are exact substrings of the generated content, following established methodologies. For ELI5, we utilize the most precise automatic metric to date, claim recall, employing the TRUE (Honovich et al., 2022) natural language inference model to ascertain if the generated output encompasses the sub-claims of the reference answer.
We present the task criteria to humans and ask them to provide a score for each answer based on the criteria. Participants are provided with a reference answer—not as an exclusive ground truth but as a guide—and are permitted to use search engines for additional context. We enlisted three annotators to evaluate 75 samples each for ASQA and ELI5, and calculate the average of them as a human score. Our analysis compares these human scores with those generated by task-specific metrics and GPT-4. The results, as detailed in our Table 4, underscore GPT-4’s closer alignment with human judgments in both ASQA and ELI5. As shown in Figure 12, as the task becomes more open-ended like Eli5, the human agreements become lower than the factoid dataset ASQA. This further evidence the assumption that the correctness of a long-form answer should be a distribution.
ASQA ELI5
Metric EM GPT-4 Claim GPT-4
Corr $\uparrow$ 47.8 76.2 42.9 71.5
MAE $\downarrow$ 43.0 12.9 52.0 9.1
Table 4: Comparison of Human, Task-Specific metric, and GPT-4 Correctness Distributions in ASQA and ELI5 Tasks. Results are with %. ‘Corr’ denotes Pearson correlation (the higher the better), ‘MAE’ denotes mean absolute error (the lower the better). ‘EM’ denotes EM-recall, and ‘Claim’ denotes Claim-recall.
Figure 6: Correlation for ASQA
Figure 7: Cohen Kappa for ASQA
Figure 8: MAE for ASQA
Figure 9: Correlation for ELI5
Figure 10: Cohen Kappa for ELI5
Figure 11: MAE for ELI5
Figure 12: Pairwise annotator agreements for ASQA and ELI5 respectively
Appendix D Main Experiment
@@ D.1 Generation
D.2 Correctness Evaluation
For ASQA, ELI5, CNNDM, we ask the GPT-4 to evaluate the correctness of examples. For ASQA and ELI5, We ask GPT-4 to evaluate the answer three times, producing a distribution capturing criteria ambiguity and model’s inherent subjectivity. However, for CNNDM, we only ask GPT-4 to evaluate once to save computation given that most answers correctness concentrates around 0.8 with small variance. The GPT-4 evaluation template can be found in Appendix F.1 and criteria for different tasks can be found in Appendix F.2. For QAMPARI, which involves generating a list of entities as answers, we determine that F1 scores provide a more suitable measure of correctness than GPT-4 evaluations. To this end, we adopt the F1-5 metric (Gao et al., 2023), computing the F1 score based on direct matches with the correct answer list and assigning a recall of 100% for responses containing at least five accurate answers (recall-5).
D.3 Confidence Elicitation
@@
Self-Evaluation We prompt the model to self-evaluate an answer 10 times, creating a confidence distribution based on its self-evaluation scores. Each self-evaluation includes task instructions, grading criteria, and evaluation examples (three for all long-form QA datasets and one for CNNDM due to length constraints). Additionally, it contains specific instructions for self-evaluation, incorporating both the question and the answer under evaluation. The detailed self-evaluation template and criteria are available in Appendix F.3 and Appendix F.2 respectively.
@@
For ASQA, we employ a self-consistency-claim approach. To assess the similarity between the two answers, we first identify factual claims in the first answer using a ClaimBuster-trained DeBERTa V2 (He et al., 2020) claim detector. We then verify the presence of these claims in the second answer through NLI. The similarity score is the average presence of factual claims across answers.
@@
For ELI5, where answers provide easily understandable explanations, we focus on the overall content. Here, we apply a self-consistency-naive method, assign a similarity score to each answer pair with simple prompting.
@@
In CNNDM, where answers are summaries highlighting key points of an article, we gauge the similarity between two answers by evaluating the overlap of key points. To achieve this, we implement the self-consistency-split method. We dissect the first answer into individual sentences and then use NLI to determine if each sentence is present in the second answer. The similarity score is derived by averaging the presence of these segmented sentences in the comparative answer.
Appendix E Improving Calibration
E.1 Fine-tuning
In our study, we focus on improving model calibration on the ASQA dataset. We explore three different strategies: fine-tune the model to do self-evaluation (Input: question and model’s answer; Output: answer’s score and explanation), fine-tune the model to do generation (Input: question; Output: answer), and a hybrid of both.
@@ E.1 Fine-tuning
Data We generate a self-evaluation training dataset with GPT-4, which evaluates different models responses to questions drawn from ASQA training set. This dataset comprises inputs of questions and corresponding model answers, with outputs including a score and an explanation for each answer. We create the self-evaluation data in two steps. The initial phase involves the creation of a diverse answer pool. This is achieved by employing a suite of models with varying computational capacities, including Llama-2-7b, Llama-2-13b, Llama2-70b, Vicuna-13b, and ChatGPT. Each model generates responses to a spectrum of questions drawn from the training set of ASQA task, ensuring the resultant answer pool encompasses a broad quality spectrum, from low to high. Subsequently, we employ GPT-4 to critically assess these answers, assigning a score and providing a corresponding explanation for its evaluation. This process yields a rich dataset, each instance of which encompasses a question, a model-generated answer, an evaluative score, and a justification for that scoring. This approach results in a comprehensive dataset with 2,000 self-evaluation examples (1,800 for training and 200 for validation), each including a question, model-generated answer, score, and justification. For generation data, we use 80% of training ASQA’s dataset (4,353 examples) for training and the remaining 20% for validation. The hybrid dataset combines the self-evaluation and generation training sets, using the self-evaluation validation set for assessment.
@@
Training and Results Regarding training, we fine-tune Llama2-13b-chat model using LORA on this dataset. We maintain consistent parameters across all scenarios: a learning rate of 5e-6, five epochs, 100 warm-up steps, and a total batch size of 4 (achieved through 4 gradient accumulation steps across four GPUs, with a batch size of 1 per device). As Table 3 reveals, solely training on self-evaluation (‘Evaluation’) did not yield consistent improvements in calibration, possibly due to the complexity of this task, as well as the limitation of LORA fine-tuning. Nonetheless, fine-tuning the model improves the self-consistency method, especially when the generation data is included during training (‘Eval + Gen’ and ‘Generation’). The model becomes more confident in terms of self-consistency after fine-tuning.
E.2 Source Documents
We investigate the effect of additional context on LLM calibration, focusing on self-consistency confidence with Llama-2-13b-chat and GPT-3.5-turbo models on the ASQA dataset. We test the model’s performance when supplemented with two different types of source documents: random documents related to the question and ‘oracle’ documents directly relevant to the answers, as included in the dataset release Stelmakh et al. (2022). Findings in Table 5 reveal that Oracle documents can enhance model performance and calibration across two out of three metrics for both models, while random documents are less effective. These results underscore the importance of relevant contextual information in model calibration.
Model Doc Corr ECE-M F10.8 Score
Llama213b N 48.0-- 15.9$\uparrow$ 3.6-- 51.3--
R 39.0 $\downarrow$ 14.8$\downarrow$ 30.6 $\uparrow$ 59.6 $\uparrow$
O 49.5 $\uparrow$ 17.5$\uparrow$ 20.1 $\uparrow$ 65.8 $\uparrow$
GPT-3.5turbo N 30.5-- 26.7$\uparrow$ 58.4-- 72.6$\downarrow$
R 36.6 $\uparrow$ 26.4$\downarrow$ 68.2 $\uparrow$ 64.8 $\downarrow$
O 27.0 $\downarrow$ 24.9$\downarrow$ 72.8 $\uparrow$ 75.8 $\uparrow$
Table 5: How source documents affect the calibration score. In the document column, “N” means no documents in the input prompt, “R” means randomly selected documents relevant to the topic, and “O” means the oracle documents relevant to the answer. $\uparrow$ denotes that the calibration score goes up when adding documents, while $\downarrow$ means going down. For ECE-M it’s opposite.
E.3 Hybrid Confidence Elicitation
We explore whether combining self-evaluation and self-consistency can yield a more accurate confidence distribution on ASQA dataset. By blending confidence distributions from self-evaluation ($C^{eval}_{i}$) and self-consistency ($C^{consist}_{i}$) into a hybrid distribution $C^{hybrid}_{i}=\alpha C^{eval}_{i}+(1-\alpha)C^{consist}_{i}$, we adjust their relative contributions using $\alpha$. As shown in Figure 13, we observe that the correlation between confidence and correctness initially increases but then declines as $\alpha$ varies from 0 to 1. However, this trend doesn’t extend to other metrics like F1, indicating that while hybrid calibration elicitation may enhance calibration in terms of correlation, it may not have the same impact on other dimensions.
@@ E.3 Hybrid Confidence Elicitation
Appendix F Prompts